OpenStudy (anonymous):

Equation of a circle in standard form Center=(0,0) R=10

OpenStudy (anonymous):

x^2+y^2=100

OpenStudy (saifoo.khan):

she cheated mine^

OpenStudy (anonymous):

No, then why did u post this question in the chat?

OpenStudy (anonymous):

so I put the center in 8^2+-6^2=100?

OpenStudy (anonymous):

I mean 0,0 I have 2 problems like this

OpenStudy (anonymous):

For my problem number one the center is (0,0) and my 2nd problem is (8,-6) both R= 10

OpenStudy (anonymous):

So then it would be (x-8)^2+(y+6)^2=100

OpenStudy (anonymous):

Here, try this site...it might help. http://www.regentsprep.org/Regents/math/algtrig/ATC1/circlelesson.htm

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No Problem :D